40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1027     Entry time: Mon Oct 6 10:00:49 2008
Author: steve 
Type: Update 
Category: MOPA 
Subject: MOPA_HTEMP is up 
Monday morning conditions:

The laser head temp is up to 20.5 C
The laser shut down on Friday without any good reason.
I was expecting the temp to come down slowly. It did not.
The control room temp is 73-74 F, Matt Evans air deflector in perfect position.
The laser chiller temp is 22.2 C

ISS is saturating. Alarm is on. Turning gain down from 7 to 2 pleases alarm handler.

c1LSC computer is down
Attachment 1: htup.jpg  88 kB  | Hide | Hide all
htup.jpg
ELOG V3.1.3-